### v1

@@ I Introduction
The stellar-mass binary black holes detectable by LIGO (LIGO Scientific Collaboration et al., 2015; Capote et al., 2025; Soni et al., 2025), Virgo (Acernese et al., 2015) and KAGRA (Akutsu et al., 2021) (LVK) are likely created from the collapse of massive stars. These stars may be born as a binary system in the galactic field that then evolves into a binary black hole system, or they may be born in a dense stellar environment such as a star cluster, in which the black hole stellar remnants dynamically assemble into tightly-bound binaries, and typically encounter other objects in the cluster before merger (see reviews by Mapelli, 2020; Mandel and Farmer, 2022, and references therein). While it is not yet known which of these potential formation scenarios, if any, contributes to the majority of binary black hole systems, one clear signature of the dynamical assembly hypothesis is the existence of hierarchical mergers (e.g. Gerosa and Fishbach, 2021). Hierarchical mergers include one or more “second-generation” remnants of previous, “first-generation” black hole mergers, and are informative for a number of reasons.
Firstly, they are rarely produced in clusters with low stellar mass, as the recoil kicks that “second-generation” black holes receive from anisotropic emission of gravitational waves are typically larger than the clusters’ escape velocity. Thus, “second-generation” black holes are often ejected from star clusters and only a small fraction remain to undergo another merger  (Fitchett, 1983; Portegies Zwart and McMillan, 2000; Favata et al., 2004; González et al., 2007; Lousto et al., 2010; Gerosa et al., 2018; Mahapatra et al., 2021). The mass and spin of the “second-generation” object in a hierarchically-merged system therefore provide a lower limit on the escape velocity of its host and insight into the system’s environment.
@@
Secondly, if assumptions about the typical star clusters that host dynamically-formed binary black holes are made, the number of mergers that include a “second-generation” black hole strongly informs the number of dynamically-formed “first-generation”+“first-generation” mergers in the population (Kimball et al., 2020), thereby constraining the total rate of dynamically assembled systems.
Finally, hierarchical mergers have clear population predictions that allow them to be straightforwardly distinguished from the larger binary black hole population (Fishbach et al., 2017; Gerosa and Berti, 2017). An ensemble of black holes that have resulted from previous mergers will have a spin magnitude distribution peaked near $\chi=0.69$ (Berti and Volonteri, 2008; Hofmann et al., 2016; Rodriguez et al., 2019; Borchers et al., 2025), regardless of how many previous mergers occurred (Fishbach et al., 2017). Additionally, “second-generation”+“first-generation” mergers will have a mass ratio distribution peaked near $\approx 1/2$ (Gerosa and Berti, 2017; Rodriguez et al., 2019; Kimball et al., 2020), and binary black holes formed in gas-free star clusters should have an isotropic spin tilt distribution (but cf. Kıroğlu et al. 2025). However, mass ratio and spin orientation distributions differ for 2G+1G mergers formed in the disks of active galactic nuclei (Tagawa et al., 2020; McKernan et al., 2020; Vaccaro et al., 2024; Vaccaro, 2025; Li et al., 2025b).
@@
In this work, we isolate the hierarchical mergers in the latest gravitational wave data using a phenomenological model that takes advantage of the most straightforward and robust population prediction for hierarchical mergers: their primary spin magnitude distribution. We then investigate the population properties of this hierarchically-merged subpopulation and provide insight into their cluster environments and the population of “first-generation” black holes in star clusters (via the 1G secondaries of 2G+1G mergers). The redshift distribution of our identified hierarchically-merged subpopulation is likely steeper than that of the rest of the population (Fig. 1). This is in contrast to naïve expectations from cluster dynamics, in which hierarchical mergers occur at similar (or slightly later) times in a cluster’s history since they necessarily occur after the first “first-generation” mergers. This finding therefore either implies that the subpopulation of $\chi_{1}\approx 0.69$ objects is not formed hierarchically, or that the hierarchical mergers are coming from a distinct population of more distant star clusters than produced our observed first-generation mergers. Regardless, the fact that a highly-spinning subpopulation has a steeper redshift evolution than the rest of the binary black holes offers a potential explanation to the previously-discovered broadening of the effective spin distribution with redshift. In a companion paper (Vijaykumar et al., in prep.), we show that this subpopulation can also explain the effective spin distribution’s narrowing with mass ratio (Callister et al., 2021; The LIGO Scientific Collaboration et al., 2025).
@@ II Phenomenological population model
Marginalized over the range $z=[0,2]$, our inferred value of $\xi=0.23^{+0.18}_{-0.17}$ is roughly consistent with previous studies that also identify a 2G+1G-like subpopulation that makes up roughly $12-20\%$ of the population [^1] Antonini et al. (2025) estimate that $20\%$ of all mergers are 2G+1G, and Tong et al. (2025a) estimate the rate of 2G+1G mergers to be $2.2^{+1.9}_{-1.2}\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, which corresponds to $\approx 12\%$ of the total BBH merger rate inferred by The LIGO Scientific Collaboration et al. (2025), albeit with a different population model. Kimball et al. (2021) find $\mathcal{R}_{\rm 2G+1G}/\mathcal{R}_{\rm 1G+1G}<0.12$ when analyzing GWTC-2, and our posterior on this value peaks at $0.23$, putting these two studies performed on different datasets in mild tension. (e.g. Kimball et al., 2021; Wang et al., 2022; Mould et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tong et al., 2025b, a; Berti et al., 2025; Vijaykumar et al., in prep.). The fact that similar fractions of 2G+1G systems are inferred – regardless of what compact binary parameter (i.e. effective spin, mass ratio, or primary spin magnitudes) is used to identify them – strengthens the evidence that these independently-discovered subpopulations are indeed the same as one another and represent 2G+1G mergers.
@@ VI Hierarchical Mergers Can Source Correlations Between Effective Spin, Redshift, Mass, and Mass Ratio
Fig. 4 summarizes the results of our analysis in the context of population-level correlations. Note that the effective spin ($\chi_{\rm eff}$) distribution cannot be analytically calculated from our component spin model, so it is reconstructed from samples and is therefore shown as a histogram, marginalized over hyperparameter uncertainty. All pairs of parameters shown in Fig. 4 exhibit correlations in the total population (orange contours), even though the correlations are absent in the individual subpopulations. In particular, the hierarchical subpopulation’s high primary spin magnitudes and isotropic spin tilts (see Appendix C) cause it to have a broader $\chi_{\rm eff}$ distribution than our main population (center panel). Its low contribution to the merger rate at low redshift but steep increase in contribution at high redshift therefore causes the overall-$\chi_{\rm eff}$ distribution to broaden with redshift (bottom center panel), offering an explanation to the correlation first discovered by Biscoveanu et al. (2022).
In a companion paper (Vijaykumar et al., in prep.), we show that a subpopulation of hierarchical mergers also offers a natural explanation of the correlation first discovered in Callister et al. (2021), and then found by The LIGO Scientific Collaboration et al. (2025) to be a narrowing of the effective spin distribution with mass ratio in GWTC-4 (but see Li et al., 2025a, for an explanation from active galactic nuclei). All previously-identified correlations in the GW data can therefore be attributed to a single origin.
@@ VII Discussion
Using the prediction that second-generation BHs should have dimensionless spin magnitudes of $\chi\approx 0.69$, we identify a subpopulation of hierarchical mergers in the fourth gravitational-wave transient catalog. The merger rate and primary mass distribution of this subpopulation is consistent with those found previously with alternative methods (Tiwari and Fairhurst, 2021; Mould et al., 2022; Wang et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tiwari, 2025; Tong et al., 2025a; Vijaykumar et al., in prep.). Having isolated this hierarchically-formed subpopulation, we study its component mass, spin orientation, and redshift distributions.
@@
The steeper redshift evolution of the 2G+1G population when compared to the 1G+1G population offers an explanation for the effective spin distribution’s broadening with redshift (Biscoveanu et al., 2022). If all observed correlations in the gravitational wave data can be attributed to a subpopulation of hierarchical mergers (see e.g. Vijaykumar et al., in prep.), a consistent picture of the binary black hole population is beginning to emerge. This allows us to begin interpreting the rest of the binary black hole population, using the 2G+1G subpopulation as an anchor. Furthermore, the inferred population properties of hierarchical mergers allows us to understand the properties of star clusters from which they are sourced.
@@ Appendix A Population model details
The mass distribution for both populations uses a pairing function formalism Fishbach and Holz (2020) with a Gaussian pairing function in mass ratio, $q=m_{2}/m_{1}$, truncated in the range $(0,1]$. Inspired by Tong et al. (2025b) and Farah et al. (2024), the hyperparameters governing both components are identical for the main population, except that we allow differing maximum masses, $m_{i}^{\max}$, for the primary and secondary components. For the subpopulation, all hyperparameters are allowed to differ between the primary and secondary masses as we expect 1G BHs to have a different mass distribution than 2G BHs. All component mass distributions are parametrized by the Broken Power Law + 2 Peaks primary mass distribution from The LIGO Scientific Collaboration et al. (2025). The resulting two-dimensional mass distribution is then
$p_{j}(m_{1},m_{2}|\lambda_{j})\propto\text{BP2P}(m_{1}|\lambda_{j}^{m_{1}})\text{BP2P}(m_{2}|\lambda_{j}^{m_{2}})\mathcal{N}(q|\mu^{q}_{j},\sigma^{q}_{j}).$ (A1)
For a description of hyperparameters governing the Broken Power Law + 2 Peaks model, we refer readers to Appendix B.3 of The LIGO Scientific Collaboration et al. (2025).

### v2

@@ I Introduction
Star clusters are foundational to much of our understanding of galactic and stellar astrophysics, as they play a large role in galactic structure and dynamics (Harris and Racine, 1979), significantly contribute to the chemical evolution of their galaxies (Shara and Hurley, 2002), and can be used to trace galaxy assembly (Brodie and Strader, 2006). Crucial but uncertain features of the star cluster population, such as its mass function and its evolution with redshift (Forbes et al., 2018), can be probed by the stellar-mass binary black holes observed by the LIGO (LIGO Scientific Collaboration et al., 2015; Capote et al., 2025; Soni et al., 2025), Virgo (Acernese et al., 2015) and KAGRA (Akutsu et al., 2021) (LVK) gravitational-wave detectors (Fishbach and Fragione, 2023). However, this requires knowing which portion of the binary black hole population is produced by clusters.
The clearest population-level signature of binary black hole formation in star clusters is the existence of hierarchical mergers (e.g. Gerosa and Fishbach, 2021). These include one or more “second-generation” remnants of previous, “first-generation” black hole mergers. The number of mergers that include a “second-generation” black hole strongly informs the total rate of cluster-origin binary black holes. In this Letter, we investigate the population properties of the hierarchically-merged binary black holes, which can be used to inform the redshift and metallicity distributions of star clusters.
Beyond their ability to identify which portion of the BBH population is sourced from clusters, hierarchical mergers are informative in their own right. Firstly, they are rarely produced in clusters with low stellar mass, as the recoil kicks that “second-generation” black holes receive from anisotropic emission of gravitational waves are typically larger than the clusters’ escape velocity. Thus, “second-generation” black holes are often ejected from star clusters and only a small fraction remain to undergo another merger  (Fitchett, 1983; Portegies Zwart and McMillan, 2000; Favata et al., 2004; González et al., 2007; Lousto et al., 2010; Gerosa et al., 2018; Mahapatra et al., 2021). The mass and spin of the “second-generation” object in a hierarchically-merged system therefore provide a lower limit on the escape velocity of its host and insight into the system’s environment.
@@
Secondly, hierarchical mergers have clear population predictions that allow them to be straightforwardly distinguished from the larger binary black hole population (Fishbach et al., 2017; Gerosa and Berti, 2017). An ensemble of black holes that have resulted from previous mergers will have a spin magnitude distribution peaked near $\chi=0.69$ (Berti and Volonteri, 2008; Hofmann et al., 2016; Rodriguez et al., 2019; Borchers et al., 2025), regardless of how many previous mergers occurred (Fishbach et al., 2017). Additionally, “second-generation”+“first-generation” mergers will have a mass ratio distribution peaked near $\approx 1/2$ (Gerosa and Berti, 2017; Rodriguez et al., 2019; Kimball et al., 2020), and binary black holes formed in gas-free star clusters should have an isotropic spin tilt distribution (but cf. Kıroğlu et al. 2025). However, mass ratio and spin orientation distributions differ for 2G+1G mergers formed in the disks of active galactic nuclei (Tagawa et al., 2020; McKernan et al., 2020; Vaccaro et al., 2024; Vaccaro, 2025; Li et al., 2025b).
@@
In this work, we isolate the hierarchical mergers in the latest gravitational wave data using a phenomenological model that takes advantage of the most straightforward and robust population prediction for hierarchical mergers: their primary spin magnitude distribution. We then investigate the population properties of this hierarchically-merged subpopulation and provide insight into their cluster environments and the population of “first-generation” black holes in star clusters (via the 1G secondaries of 2G+1G mergers). The redshift distribution of our identified hierarchically-merged subpopulation is likely steeper than that of the rest of the population (Fig. 1). This is in contrast to naïve expectations from cluster dynamics, in which hierarchical mergers occur at similar (or slightly later) times in a cluster’s history since they necessarily occur after the first “first-generation” mergers. This finding therefore either implies that the subpopulation of $\chi_{1}\approx 0.69$ objects is not formed hierarchically, or that the hierarchical mergers are coming from a distinct population of more distant star clusters than produced our observed first-generation mergers. Regardless, the fact that a highly-spinning subpopulation has a steeper redshift evolution than the rest of the binary black holes offers a potential explanation to the previously-discovered broadening of the effective spin distribution with redshift. In a companion paper (Vijaykumar et al., 2026), we show that this subpopulation can also explain the effective spin distribution’s narrowing with mass ratio (Callister et al., 2021; The LIGO Scientific Collaboration et al., 2025).
@@ II Phenomenological population model
Marginalized over the range $z=[0,2]$, our inferred value of $\xi=0.23^{+0.18}_{-0.17}$ is roughly consistent with previous studies that also identify a 2G+1G-like subpopulation that makes up roughly $12-20\%$ of the population [^1] Antonini et al. (2025) estimate that $20\%$ of all mergers are 2G+1G, and Tong et al. (2025a) estimate the rate of 2G+1G mergers to be $2.2^{+1.9}_{-1.2}\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, which corresponds to $\approx 12\%$ of the total BBH merger rate inferred by The LIGO Scientific Collaboration et al. (2025), albeit with a different population model. Pierra et al. (2024) find that $\approx 2$$5\%$ of the overall population belong to a $\chi\approx 0.7$ subpopulation, which is lower than that inferred in this work, potentially caused by differing inferred redshift distributions. Additionally, Kimball et al. (2021) find $\mathcal{R}_{\rm 2G+1G}/\mathcal{R}_{\rm 1G+1G}<0.12$ when analyzing GWTC-2, and our posterior on this value peaks at $0.23$, putting these two studies performed on different datasets in mild tension. (e.g. Kimball et al., 2021; Wang et al., 2022; Mould et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tong et al., 2025b, a; Berti et al., 2025; Vijaykumar et al., 2026). The fact that similar fractions of 2G+1G systems are inferred – regardless of what compact binary parameter (i.e. effective spin, mass ratio, or primary spin magnitudes) is used to identify them – strengthens the evidence that these independently-discovered subpopulations are indeed the same as one another and represent 2G+1G mergers.
@@ VI Hierarchical Mergers Can Source Correlations Between Effective Spin, Redshift, Mass, and Mass Ratio
Fig. 4 summarizes the results of our analysis in the context of population-level correlations. Note that the effective spin ($\chi_{\rm eff}$) distribution cannot be analytically calculated from our component spin model, so it is reconstructed from samples and is therefore shown as a histogram, marginalized over hyperparameter uncertainty. Additionally, the redshift distribution displayed in Fig. 4 (bottom-right panel) is the (normalized) probability distribution over redshift, $p(z)$, – as opposed to the volumetric merger rate shown in Fig. 1 – because $p(z)$ is directly obtained by marginalizing the two-dimensional probability distributions shown in the off-diagonal panels.
All pairs of parameters shown in Fig. 4 exhibit correlations in the total population (orange contours), even though the correlations are absent in the individual subpopulations. In particular, the hierarchical subpopulation’s high primary spin magnitudes and isotropic spin tilts (see Appendix C) cause it to have a broader $\chi_{\rm eff}$ distribution than our main population (center panel). Its low contribution to the merger rate at low redshift but steep increase in contribution at high redshift therefore causes the overall-$\chi_{\rm eff}$ distribution to broaden with redshift (bottom center panel), offering an explanation to the correlation first discovered by Biscoveanu et al. (2022).
In a companion paper (Vijaykumar et al., 2026), we show that a subpopulation of hierarchical mergers also offers a natural explanation of the correlation first discovered in Callister et al. (2021), and then found by The LIGO Scientific Collaboration et al. (2025) to be a narrowing of the effective spin distribution with mass ratio in GWTC-4 (but see Li et al., 2025a, for an explanation from active galactic nuclei). All previously-identified correlations in the GW data can therefore be attributed to a single origin.
@@ VII Discussion
Using the prediction that second-generation BHs should have dimensionless spin magnitudes of $\chi\approx 0.69$, we identify a subpopulation of hierarchical mergers in the fourth gravitational-wave transient catalog. The merger rate and primary mass distribution of this subpopulation is consistent with those found previously with alternative methods (Tiwari and Fairhurst, 2021; Mould et al., 2022; Wang et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tiwari, 2025; Tong et al., 2025a; Vijaykumar et al., 2026). Having isolated this hierarchically-formed subpopulation, we study its component mass, spin orientation, and redshift distributions.
@@
The steeper redshift evolution of the 2G+1G population when compared to the 1G+1G population offers an explanation for the effective spin distribution’s broadening with redshift (Biscoveanu et al., 2022). If all observed correlations in the gravitational wave data can be attributed to a subpopulation of hierarchical mergers (see e.g. Vijaykumar et al., 2026), a consistent picture of the binary black hole population is beginning to emerge. This allows us to begin interpreting the rest of the binary black hole population, using the 2G+1G subpopulation as an anchor.
The population properties of hierarchical mergers inferred in this work will enable constraints on the properties of star clusters from which they are sourced. For instance, interpreting the secondary mass distribution of the 2G+1G subpopulation as the birth mass distribution of BHs in star clusters can inform the metallicity content of these clusters, as well as the supernova mechanism. Additionally, the the 2G+1G subpopulation’s redshift distribution informs the formation times, metallicity content, and redshift-dependent mass function of star clusters.
@@ Appendix A Population model details
The mass distribution for both populations uses a pairing function formalism (Fishbach and Holz, 2020) with a Gaussian pairing function in mass ratio, $q=m_{2}/m_{1}$, truncated in the range $(0,1]$. Inspired by Tong et al. (2025b) and Farah et al. (2024), the hyperparameters governing both components are identical for the main population, except that we allow differing maximum masses for the primary and secondary components. For the subpopulation, all hyperparameters are allowed to differ between the primary and secondary masses, as we expect 1G BHs to have a different mass distribution than 2G BHs. All component mass distributions are parametrized by the Broken Power Law + 2 Peaks primary mass distribution from The LIGO Scientific Collaboration et al. (2025). For a given subpopulation, the resulting two-dimensional mass distribution is then
$p_{j}(m_{1},m_{2}|\lambda_{j})\propto\text{BP2P}(m_{1}|\lambda_{j}^{m_{1}})\text{BP2P}(m_{2}|\lambda_{j}^{m_{2}})\mathcal{N}(q|\mu^{q}_{j},\sigma^{q}_{j}),$ (A1)
where $j=[S,M]$ denotes the particular subpopulation. For a description of hyperparameters governing the Broken Power Law + 2 Peaks model, we refer readers to Appendix B.3 of The LIGO Scientific Collaboration et al. (2025).
@@
Both subpopulations have redshift distributions of the form (Fishbach et al., 2018)
$p_{j}(z|\kappa_{j})\propto\frac{{\rm d}V_{c}}{{\rm d}z}(1+z)^{\kappa_{j}-1}.$ (A2)